# From Bugs to Benchmarks: A Comprehensive Survey of Software Defect Datasets

### v1

@@ 3.2.1. Functional Defects
Functional bugs are the most common focus in software defect datasets, as they form the backbone of software reliability studies. These datasets either provide a general collection of functional bugs across various categories (e.g., Defects4J (Just et al., 2014), Bears (Madeiral et al., 2019), Bugs.jar (Saha et al., 2018), BugSwarm (Tomassi et al., 2019)) or target a specific subset of functional bugs to facilitate detailed analysis and testing. Examples of such specialized datasets include API misuses (Wang et al., 2017), regular expression defects (Wang et al., 2022a), single-statement bugs, (Karampatsis and Sutton, 2020), type errors (Zhang et al., 2022b), string related defects (Eghbali and Pradel, 2020), and numerical bugs (Wang et al., 2022b; Franco et al., 2017). Overall, functional bug datasets cover a wide variety of topics, making them versatile resources for software defect research.
@@ 3.2.2. Security Defects
Different vulnerability datasets focus on different kinds of defects. One main type of security software defect dataset, e.g., ReposVul (Wang et al., 2024), Big-Vul(Fan et al., 2020), and CVEjoin (da Ponte et al., 2023), focuses on well-defined vulnerabilities, such as those cataloged in standardized databases like CVE (Common Vulnerabilities and Exposures) and categorized by CWE (Common Weakness Enumeration). Another kind of dataset focuses on buffer overflows (Ku et al., 2007), which are critical flaws where programs write memory locations beyond buffer boundaries, and compiler-introduced security bugs (Xu et al., 2023), which arise due to incorrect compiler optimizations. Lastly, cross-language vulnerabilities, which involve multiple programming languages, are captured in datasets like CrossVul (Nikitopoulos et al., 2021), providing a basis for multi-language defense techniques.
@@ 3.2.3. Performance Defects
Performance defects are software issues that cause performance degradation, leading to inefficient use of resources, increased latency, and longer execution times. These bugs are critical in high-performance computing, web applications, and real-time systems where responsiveness and resource management are essential (Jin et al., 2012). Performance bugs can arise from improper algorithms, inefficient memory usage, excessive computation, or mismanagement of concurrency. These defects can significantly impact user experience and operational costs, making their identification and resolution a priority in software engineering. As shown in Figure 3, datasets of performance issues are prevalent in 9 of 11 application domains, with the notable exceptions of mobile applications and computational notebooks.
@@
Software defect datasets for performance issues have been constructed for both general applications and specific domains. For example, TANDEM (Sánchez et al., 2020) offers a taxonomy and dataset of real-world performance bugs, while ECench (Kim et al., 2022) introduces a benchmark of energy bugs specifically in Ethereum client software, highlighting issues where performance inefficiencies lead to excessive energy consumption. There are also datasets constructed from empirical research of performance issues, such as studies on performance degradation in browser environments (Zaman et al., 2012), in JavaScript projects (Selakovic and Pradel, 2017), in database systems (Cui et al., 2024) and applications (Liu et al., 2024b), in high-performance computing (HPC) systems (Azad et al., 2023), in numerical libraries (Franco et al., 2017), and in cloud systems (Gunawi et al., 2014). Such datasets could facilitate future research on pinpointing or mitigating performance issues.
3.2.4. Concurrency Defects
Concurrency defects arise in software when multiple threads or processes execute simultaneously and interact in unintended ways, leading to nondeterministic behavior, crashes, or data corruption. These defects are particularly challenging because they often do not manifest consistently, making them difficult to detect, reproduce, and fix. Concurrency bugs include race conditions, deadlocks, and atomicity violations, which occur when proper synchronization mechanisms fail, or incorrect access to shared resources occurs. Given the complexity of multithreaded systems and the rise of concurrent programming models, identifying and studying concurrency bugs has become critical in ensuring the stability and reliability of modern software systems. As shown in Figure 3, concurrency defects exist in application domains where parallelism and real-time execution are critical, e.g., machine learning, compilers and runtime engines, autonomous systems, and cloud computing.
Several datasets have been developed to address different aspects of concurrency bugs. For C/C++, RaceBench (Liang et al., 2023) provides triggerable and observable race condition bugs to test race detectors, and RADBench (Jalbert et al., 2011) offers a concurrency bug benchmark suite focused on race conditions, deadlocks, and atomicity violations. For Java, Jbench (Gao et al., 2018) presents a suite of data race bugs specifically designed for Java concurrency testing, and JaConTeBe (Lin et al., 2015) focuses on concurrency bugs in real-world Java projects, offering manually written test cases for verification. Additionally, there are some datasets of concurrency bugs constructed from empirical studies, e.g., on actor-based concurrency systems (Bagherzadeh et al., 2020), cloud computing systems (Gunawi et al., 2014), deep learning compilers (Shen et al., 2021), and Node.js (Wang et al., 2017), which offer invaluable insights for future studies on concurrency bugs.
3.2.5. Other Types of Defects
Some datasets do not fall neatly into the four categories of functional, security, performance, or concurrency bugs. For example, such datasets consider critical bugs that prevent other bugs from being resolved (Ren et al., 2020), deployment faults in deep learning-based mobile applications (Chen et al., 2021b), or resource leak bugs in Android applications (Liu et al., 2019). These datasets address specialized defect types that fall outside the four main categories.
3.3. Programming Languages of Defective Software
@@
Java (Andrade et al., 2024; Liu et al., 2024a; Tebib et al., 2024; Liu et al., 2024b; Xu et al., 2024; Das et al., 2024; Tian et al., 2024; Silva et al., 2024; Wang et al., 2024; Song et al., 2023; Zhong et al., 2023; Lee et al., 2024; Wu et al., 2024; Avula et al., 2023; Senanayake et al., 2023; Jiang et al., 2023; Xiong et al., 2023; Pachouly et al., 2022; Keshavarz and Nagappan, 2022; Kim et al., 2022; Bui et al., 2022; Song et al., 2022; Wendland et al., 2021; Ngo et al., 2021; Afrose et al., 2023; Makhshari and Mesbah, 2021; Ferenc et al., 2020; Karampatsis and Sutton, 2020; Bures et al., 2020; Bagherzadeh et al., 2020; Wang et al., 2020; Tomassi et al., 2019; Radu and Nadi, 2019; Riganelli et al., 2019; Madeiral et al., 2019; Saha et al., 2018; Ferenc et al., 2018; Gao et al., 2018; Madeyski and Kawalerowicz, 2017; Mitra and Ranganath, 2017; Liu et al., 2019; Lin et al., 2015; Just et al., 2014; D’Ambros et al., 2012; Chen et al., 2021b; Gunawi et al., 2014; Lin et al., 2017; Tan et al., 2018) 48
C/C++ (Timperley et al., 2024; Andrade et al., 2024; Liu et al., 2024a; Tebib et al., 2024; Zhang et al., 2024a; Jiang et al., 2024; Cui et al., 2024; Yu et al., 2024; Tian et al., 2024; Wang et al., 2024; Zhong et al., 2023; An et al., 2023; Avula et al., 2023; Kim and Hong, 2023; Liang et al., 2023; Chen et al., 2023b; Guan et al., 2023; Zhang et al., 2022a; Pereira et al., 2022; Paltenghi and Pradel, 2022; Liu et al., 2022; Kim et al., 2021b; Paul et al., 2021; Zheng et al., 2021; Du et al., 2021; Makhshari and Mesbah, 2021; Fan et al., 2020; Garcia et al., 2020; Lin et al., 2019; Tan et al., 2017; Le Goues et al., 2015, 2015; Zaman et al., 2012; Jalbert et al., 2011; Ku et al., 2007; Ho et al., 2023; Xu et al., 2023; Wang et al., 2023a; Azad et al., 2023; Zhao et al., 2023b; Wang et al., 2023c; Shen et al., 2021; Wang et al., 2021a; Gu et al., 2019; Franco et al., 2017; Lu et al., 2005; Cifuentes et al., 2009) 47
Python (Timperley et al., 2024; Liu et al., 2024a; Tian et al., 2024; Jimenez et al., 2024; Wang et al., 2024; Zhong et al., 2023; Wu et al., 2024; Avula et al., 2023; Mahbub et al., 2023; Zhao et al., 2023a; Morovati et al., 2023; Guan et al., 2023; Liang et al., 2022; Zhang et al., 2022b; Paltenghi and Pradel, 2022; Liu et al., 2022; Akimova et al., 2021; Kim et al., 2021b; Kamienski et al., 2021; Du et al., 2021; Makhshari and Mesbah, 2021; Widyasari et al., 2020; Wang et al., 2020; Tomassi et al., 2019; Radu and Nadi, 2019; Zhang et al., 2018; Chen et al., 2021b; Ho et al., 2023; de Santana et al., 2024; Zhao et al., 2023b; Wang et al., 2022b; Shen et al., 2021; Franco et al., 2017; Lin et al., 2017) 34
JavaScript (Zhong et al., 2023; Bhuiyan et al., 2023; Csuvik and Vidács, 2022; Gyimesi et al., 2021; Makhshari and Mesbah, 2021; Wang et al., 2020; Selakovic and Pradel, 2017; Wang et al., 2023a, c; Eghbali and Pradel, 2020; Wang et al., 2017) 11
Go (Cui et al., 2024; Yu et al., 2024; Xu et al., 2024; Saavedra et al., 2024; Kim et al., 2022; Yuan et al., 2021) 6
C# (Andrade et al., 2024; Zhong et al., 2023; Kim et al., 2022; Paltenghi and Pradel, 2022) 4
Rust (Zhang et al., 2024a; Yu and Wang, 2024; Kim et al., 2022; Wang et al., 2023c) 4
Solidity (Cai et al., 2024; Chen et al., 2022; Zhang et al., 2020; Wang et al., 2023b) 4
Kotlin (Tebib et al., 2024; Xiong et al., 2023; Benton et al., 2019) 3
PHP (Andrade et al., 2024; Pramod et al., 2024) 2
Others Fortran (Franco et al., 2017), F# (Paltenghi and Pradel, 2022), Groovy (Benton et al., 2019), Haskell (Applis and Panichella, 2023), Julia (Murphy et al., 2020), Q# (Paltenghi and Pradel, 2022), Ruby (Madeyski and Kawalerowicz, 2017), Scala (Bagherzadeh et al., 2020) 8
Another way of categorizing datasets is by the programming language in which the defective code is written. 100/132 of our surveyed defect datasets focus on one or multiple programming languages. These languages include Java, Python, C/C++, JavaScript, Go, and others. Table 4 presents the distribution of programming languages across software defect datasets. Java stands out as the most represented language, with a total of 48 (48/132) datasets, followed by C/C++ (47 datasets, 47/132) and Python (34 datasets, 34/132). The dominance of these languages is likely due to their popularity in software development and due to the focus of the research community on these languages. Other languages, such as F#, Groovy, Haskell, and Julia, have fewer datasets available.
The distribution of languages targeted in defect datasets has evolved over time. Figure 4 highlights the trend for the top five programming languages in software defect datasets: Java, C/C++, Python, JavaScript, and Go. The number of datasets related to Java has seen a steady increase over the past ten years, while Python has gained significant popularity since 2017. JavaScript and Go have experienced their own peaks within the past five years. C/C++ datasets have consistently appeared over time. It is important to note the surge in the number of software defect datasets since 2023. While the general increase in publication volume contributes to this rise, another factor could be the growing demand for training datasets, e.g., to train or fine-tune large language models. This trend will be further discussed in Section 6.

### v2

@@ 3.2.1. Functional Defects
Functional bugs are the most common focus in software defect datasets, as they form the backbone of software reliability studies. These datasets either provide a general collection of functional bugs across various categories (e.g., Defects4J (DBLP:conf/issta/JustJE14), Bears (DBLP:conf/wcre/DelfimUMM19), Bugs.jar (DBLP:conf/msr/SahaLLYP18), BugSwarm (DBLP:conf/icse/DmeiriTWBLDVR19)) or target a specific subset of functional bugs to facilitate detailed analysis and testing. Examples of such specialized datasets include API misuses (DBLP:conf/kbse/WangDGGQYW17), regular expression defects (Wang et al., 2022), single-statement bugs, (DBLP:conf/msr/KarampatsisS20), type errors (DBLP:conf/wcre/ZhangYYCYZ22), string related defects (DBLP:conf/kbse/EghbaliP20), and numerical bugs (DBLP:conf/kbse/WangWCCY22; DBLP:conf/kbse/FrancoGR17). Overall, functional bug datasets cover a wide variety of topics, making them versatile resources for software defect research.
@@ 3.2.2. Security Defects
Different vulnerability datasets focus on different kinds of defects. One main type of security software defect dataset, e.g., ReposVul (DBLP:conf/icse/WangHGWC024), Big-Vul(DBLP:conf/msr/FanL0N20), and CVEjoin (DBLP:conf/aina/PonteRM23a), focuses on well-defined vulnerabilities, such as those cataloged in standardized databases like CVE (Common Vulnerabilities and Exposures) and categorized by CWE (Common Weakness Enumeration). Another kind of dataset focuses on buffer overflows (DBLP:conf/kbse/KuHCL07), which are critical flaws where programs write memory locations beyond buffer boundaries, and compiler-introduced security bugs (DBLP:conf/uss/XuLDDLWPM23), which arise due to incorrect compiler optimizations. Lastly, cross-language vulnerabilities, which involve multiple programming languages, are captured in datasets like CrossVul (DBLP:conf/sigsoft/NikitopoulosDLM21), providing a basis for multi-language defense techniques.
@@ 3.2.3. Performance Defects
Performance defects are software issues that cause performance degradation, leading to inefficient use of resources, increased latency, and longer execution times. These bugs are critical in high-performance computing, web applications, and real-time systems where responsiveness and resource management are essential (Jin et al., 2012). Performance bugs can arise from improper algorithms, inefficient memory usage, excessive computation, or mismanagement of concurrency. These defects can significantly impact user experience and operational costs, making their identification and resolution a priority in software engineering. As shown in Figure 4, datasets of performance issues are prevalent in 9 of 11 application domains, with the exceptions of mobile applications and computational notebooks.
@@
Software defect datasets for performance issues have been constructed for both general applications and specific domains. For example, TANDEM (DBLP:journals/access/SanchezDMS20) offers a taxonomy and dataset of real-world performance bugs, while ECench (DBLP:conf/msr/KimKL22) introduces a benchmark of energy bugs specifically in Ethereum client software, highlighting issues where performance inefficiencies lead to excessive energy consumption. There are also datasets constructed from empirical research of performance issues, such as studies on performance degradation in browser environments (DBLP:conf/msr/ZamanAH12), in JavaScript projects (DBLP:conf/se/SelakovicP17), in database systems (DBLP:conf/icse/CuiD0WSZWYXH0024) and applications (DBLP:journals/tosem/LiuMC24), in high-performance computing (HPC) systems (DBLP:conf/msr/AzadIHR23), in numerical libraries (DBLP:conf/kbse/FrancoGR17), and in cloud systems (DBLP:conf/cloud/GunawiHLPDAELLM14). Such datasets could facilitate future research on pinpointing or mitigating performance issues.
@@
Java (DBLP:conf/edcc/AndradeLV24; DBLP:conf/msr/LiuHLZCSHM24; DBLP:conf/pst/TebibAAG24; DBLP:journals/tosem/LiuMC24; DBLP:conf/issta/XuG024; DBLP:conf/kbse/DasAM24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/msr/SilvaSM24; DBLP:conf/icse/WangHGWC024; DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/icst/LeeKYY24; DBLP:conf/sigsoft/WuLZ024; DBLP:conf/kbse/AvulaVM23; DBLP:conf/secrypt/SenanayakeKAP023; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:journals/softx/PachoulyAK22; DBLP:conf/msr/KeshavarzN22; DBLP:conf/msr/KimKL22; DBLP:conf/msr/BuiSF22; DBLP:conf/issta/Song0NW0DM22; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:conf/splc/NgoNNV21; DBLP:journals/tse/AfroseXRMY23; DBLP:conf/icse/Makhshari021; DBLP:journals/jss/FerencGGTG20; DBLP:conf/msr/KarampatsisS20; DBLP:conf/icst/BuresHA20; DBLP:journals/pacmpl/0001FSK20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/msr/RaduN19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/msr/SahaLLYP18; DBLP:conf/promise/FerencTLSG18; DBLP:conf/msr/GaoYJLYZ18; DBLP:conf/msr/MadeyskiK17; DBLP:conf/promise/MitraR17; DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/kbse/LinMZCZ15; DBLP:conf/issta/JustJE14; DBLP:journals/ese/DAmbrosLR12; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/oopsla/LinKCS17; DBLP:conf/icse/TanDGR18; DBLP:conf/fie/SandersWA24; DBLP:conf/issta/WuJPLD0BS23; DBLP:conf/msr/AmannNNNM16; DBLP:conf/sigsoft/GaoDQGW0HZW18; DBLP:conf/sigsoft/Zhang0C0Z19; DBLP:conf/wcre/ReyesGSBM24; DBLP:journals/tosem/TufanoWBPWP19; DBLP:journals/tse/ZhaoXBCL23; DBLP:conf/kbse/XuZL23) 57
C/C++ (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/edcc/AndradeLV24; DBLP:conf/msr/LiuHLZCSHM24; DBLP:conf/pst/TebibAAG24; DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:journals/tosem/JiangJWMLZ24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/icse/WangHGWC024; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/kbse/AnKCYY23; DBLP:conf/kbse/AvulaVM23; DBLP:conf/icst/KimH23; DBLP:conf/asiaccs/LiangYDMHZ23; DBLP:conf/raid/0001DACW23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/uss/ZhangP0W22; DBLP:conf/prdc/PereiraAV22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/msr/KimKL21; DBLP:conf/icse/PaulTB21a; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:conf/issre/Du0MZ21; DBLP:conf/icse/Makhshari021; DBLP:conf/msr/FanL0N20; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/icics/LinXZX19; DBLP:conf/icse/TanYYMR17; DBLP:journals/tse/GouesHSBDFW15; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/msr/ZamanAH12; DBLP:conf/hotpar/JalbertPPS11; DBLP:conf/kbse/KuHCL07; DBLP:conf/icsm/HoMI0SKNR23; DBLP:conf/uss/XuLDDLWPM23; DBLP:journals/infsof/WangBWYGS23; DBLP:conf/msr/AzadIHR23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/sigsoft/WangLX0S21; DBLP:conf/compsac/GuWL0019; DBLP:conf/kbse/FrancoGR17; lu2005bugbench; BegBunch; DBLP:conf/icse-apr/YadavW24; DBLP:conf/msr/BeyerGKRR24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/tse/ZhaoXBCL23; DBLP:conf/sp/Dolan-GavittHKL16) 52
Python (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/msr/LiuHLZCSHM24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/iclr/JimenezYWYPPN24; DBLP:conf/icse/WangHGWC024; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/sigsoft/WuLZ024; DBLP:conf/kbse/AvulaVM23; DBLP:conf/msr/MahbubSR23; DBLP:journals/jss/ZhaoMLZ23; DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/apsec/AkimovaBDKKMM21; DBLP:conf/msr/KimKL21; DBLP:conf/msr/KamienskiPBH21; DBLP:conf/issre/Du0MZ21; DBLP:conf/icse/Makhshari021; DBLP:conf/sigsoft/WidyasariSLQPTT20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/msr/RaduN19; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/icsm/HoMI0SKNR23; DBLP:journals/tosem/SantanaNAA24; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/kbse/WangWCCY22; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/kbse/FrancoGR17; DBLP:conf/oopsla/LinKCS17; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:journals/access/AntalVKMHF24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tse/ZhaoXBCL23) 39
JavaScript (DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/icse/BhuiyanPVPS23; DBLP:conf/msr/CsuvikV22; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/icse/Makhshari021; DBLP:conf/msr/WangBJS20; DBLP:conf/se/SelakovicP17; DBLP:journals/infsof/WangBWYGS23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/kbse/EghbaliP20; DBLP:conf/kbse/WangDGGQYW17; DBLP:conf/msr/OliverDAH24; DBLP:conf/kbse/RomanoLK021) 13
Go (DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/issta/XuG024; DBLP:conf/icse/SaavedraSM24; DBLP:conf/msr/KimKL22; DBLP:conf/cgo/YuanLLLLX21; DBLP:conf/icse/HuZLYH024) 7
C# (DBLP:conf/edcc/AndradeLV24; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/msr/KimKL22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/fie/SandersWA24) 5
Rust (DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:conf/qrs/YuW24; DBLP:conf/msr/KimKL22; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/kbse/RomanoLK021) 5
Solidity (DBLP:journals/jss/CaiLZZS24; DBLP:journals/tse/ChenXLGLC22; DBLP:conf/icsm/ZhangXL20; DBLP:journals/jss/WangCHZBZ23) 4
Kotlin (DBLP:conf/pst/TebibAAG24; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/icse/BentonGZ19) 3
PHP (DBLP:conf/edcc/AndradeLV24; DBLP:conf/msr/PramodSTSW24) 2
Ruby (DBLP:conf/msr/MadeyskiK17; DBLP:journals/pacmpl/DrososSAM024) 2
Others Fortran (DBLP:conf/kbse/FrancoGR17), F# (DBLP:journals/pacmpl/PaltenghiP22), Groovy (DBLP:conf/icse/BentonGZ19), Haskell (DBLP:conf/msr/ApplisP23), Julia (DBLP:conf/hotsos/MurphyBSR20), Q# (DBLP:journals/pacmpl/PaltenghiP22), Scala (DBLP:journals/pacmpl/0001FSK20) 7
3.2.4. Concurrency Defects
Concurrency defects arise in software when multiple threads or processes execute simultaneously and interact in unintended ways, leading to nondeterministic behavior, crashes, or data corruption. These defects are particularly challenging because they often do not manifest consistently, making them difficult to detect, reproduce, and fix. Concurrency bugs include race conditions, deadlocks, and atomicity violations, which occur when proper synchronization mechanisms fail, or incorrect access to shared resources occurs. Given the complexity of multithreaded systems and the rise of concurrent programming models, identifying and studying concurrency bugs has become critical in ensuring the stability and reliability of modern software systems. As shown in Figure 4, concurrency defects exist in application domains where parallelism and real-time execution are critical, e.g., machine learning, compilers and runtime engines, cloud computing, and autonomous systems.
Several datasets have been developed to address different aspects of concurrency bugs. For C/C++, RaceBench (DBLP:conf/asiaccs/LiangYDMHZ23) provides triggerable and observable race condition bugs to test race detectors, and RADBench (DBLP:conf/hotpar/JalbertPPS11) offers a concurrency bug benchmark suite focused on race conditions, deadlocks, and atomicity violations. For Java, Jbench (DBLP:conf/msr/GaoYJLYZ18) presents a suite of data race bugs specifically designed for concurrency testing, and JaConTeBe (DBLP:conf/kbse/LinMZCZ15) focuses on concurrency bugs in real-world projects, offering manually written test cases for verification. Additionally, there are some datasets of concurrency bugs constructed from empirical studies, e.g., on actor-based concurrency systems (DBLP:journals/pacmpl/0001FSK20), cloud computing systems (DBLP:conf/cloud/GunawiHLPDAELLM14), deep learning compilers (DBLP:conf/sigsoft/ShenM0TCC21), and Node.js (DBLP:conf/kbse/WangDGGQYW17), which offer invaluable insights for future studies.
3.2.5. Other Types of Defects
Some datasets do not fall neatly into the four categories of functional, security, performance, or concurrency bugs. For example, such datasets consider critical bugs that prevent other bugs from being resolved (DBLP:conf/iwpc/RenL020), deployment faults in deep learning-based mobile applications (DBLP:conf/icse/ChenYLCLWL21), or resource leak bugs in Android applications (DBLP:journals/ese/LiuWWXCWYZ19). These datasets address specialized defect types that fall outside the four main categories.